### v1

# Mechanistic Modeling of Continuous Lyophilization for Pharmaceutical Manufacturing

Abstract
Lyophilization (also known as freeze drying) is a process that is commonly used to increase the stability of drug products, e.g., mRNA vaccines, in pharmaceutical manufacturing. While extensive efforts have been dedicated to shift the pharmaceutical industry towards continuous manufacturing, the majority of industrial-scale lyophilization is still being operated in a batch mode. This article proposes the first mechanistic model for a complete continuous lyophilization process, which includes freezing, primary drying, and secondary drying. The state-of-the-art lyophilization technology is considered, in which vials are suspended and moved continuously through the process. The model can describe the evolution of several critical process parameters, namely the product temperature, ice/water fraction, sublimation front position, and concentration of bound water, for the entire lyophilization process. The model is also demonstrated for several applications related to process design and optimization. Ultimately, the framework and results presented in this work can serve as a solid foundation to guide the design and development of future continuous lyophilization processes.
**footnotetext: Corresponding author. Email: braatz@mit.edu
Keywords— Lyophilization, Freeze Drying, Continuous Manufacturing, Continuous Lyophilization, Pharmaceutical Manufacturing, Suspended Vials
Introduction
Lyophilization (also known as freeze drying) is a process that is commonly used to increase the stability of drug products in pharmaceutical manufacturing [1]. Various promising applications of lyophilization have been proven, including its recent applications that provide long-term stability for mRNA vaccines (e.g., for COVID-19) to allow for higher storage temperature [2, 3]. Advancement in lyophilization technology could therefore play a crucial role in the future of the pharmaceutical industry. A typical lyophilization process consists of three main steps, namely (1) freezing, (2) primary drying, and (3) secondary drying. During freezing, the product and liquid (usually water) in a vial are cooled such that most of the liquid (free water) is frozen, with a small amount of liquid (bound water) retaining its liquid state adsorbed to the solid material between the ice crystals. In primary drying, the free water in the form of ice crystals is removed via sublimation under vacuum. Finally, in secondary drying, the product is heated further such that the bound water can be removed via desorption.
Although the current trends in the pharmaceutical industry largely focus on the adoption of continuous manufacturing, the majority of production-scale lyophilization is still being operated in a batch mode [4, 5], with most research efforts dedicated to process optimization, monitoring, and control to ensure that the final product quality meets the regulations [6, 5]. Conventional lyophilization of unit doses usually entails a batch of vials containing drug products situated on the cooling/heating shelf where the entire lyophilization process occurs. A number of continuous lyophilization concepts for pharmaceutical products have been proposed, including spray freeze drying [7] and spin freezing [4] (see a comprehensive review of continuous lyophilization technologies in Ref. [5]). A novel lyophilization technology has been recently proposed by Ref. [6], in which vials are suspended and continuously moved along the process without any complicated motions (e.g., as in spin or spray freeze drying), allowing the product quality control to be done more conveniently and rigorously.
Mathematical models have been widely used to assist the design, optimization, and control of lyophilization processes. Since the phenomena in lyophilization are underpinned by several theories in the area of heat and mass transfer, models for lyophilization are mostly physics-based rather than data-driven. Mechanistic models for batch/conventional lyophilization have been extensively discussed in the literature for many decades, e.g., for freezing [8, 9, 10, 11, 12], primary drying [13, 14, 15, 16, 17, 8, 18, 19, 20, 21, 22], and secondary drying [13, 14, 15, 16, 17, 23, 21, 24, 25]. Only a few models for continuous lyophilization are available. For example, a model for the primary drying step in continuous drying of spin frozen vials was proposed by Ref. [26]. A model for the freezing step was also developed for spin freezing [27]. For spray freeze drying, Ref. [28] developed a detailed model that can predict the temperature of droplets during the cooling and freezing phases. Nevertheless, there is no model for the suspended-vial configuration. Besides, none of the published models for continuous lyophilization considers a complete lyophilization cycle (all three steps), which is critical for optimization and control of the entire continuous operation.
This article presents the first mechanistic model for continuous lyophilization of suspended vials. The model is developed to capture important phenomena of the entire lyophilization process, including cooling and stochastic/controlled ice nucleation during the freezing step, heat and mass transfer with sublimation during the primary drying step, and heat transfer with desorption during the secondary drying step. The proposed model is validated and employed to predict the evolution of important process parameters, with several applications associated with process design and optimization demonstrated.
This article is organized as follows. Section 2 gives an overview of conventional (batch) and continuous lyophilization technologies. Section 3 extensively describes the development of our mechanistic model and its implementation, supported by a comprehensive review on a variety of modeling strategies used in the literature. Section 4 validates the proposed model with the experimental data from the relevant literature. Section 5 showcases four case studies to demonstrate various applications of our model. Finally, Section 6 summarizes the study and briefly discusses possible future work.
Process Description
This work considers lyophilization of unit doses, in which the product is introduced into a vial prior to being lyophilized. This type of lyophilization, compared to lyophilization of bulk material, is more preferable in (bio)pharmaceutical applications due to its accurate dosage and better control of sterility [5]. Therefore, results and discussion presented in this article are based entirely on lyophilization of unit doses.
Batch lyophilization
The majority of lyophilization processes in the pharmaceutical industry has been operated in a batch mode for many decades. The conventional configuration typically comprises a large number of vials located on the cooling/heating shelf (Fig. 1A). The bottom shelf, whose temperature can be manipulated, is used to cool the vials during the freezing step and heat the vials during the drying steps.
Figure 1: (A) Conventional batch lyophilization of unit doses. A number of vials are placed on the cooling/heating shelf. (B) Continuous lyophilization of suspended vials. A number of vials are suspended and moved continuously along the lyophilizer.
Several drawbacks associated with batch lyophilization are extensively discussed in Refs. [6, 5]. Typical issues of any batch process include batch-to-batch variability, quality control, process downtime, and operational flexibility. Another disadvantage that is specific to the configuration illustrated in Fig. 1A is non-uniformity in heat transfer. For example, vials located on the side and corner of the shelf are typically affected by thermal radiation from the chamber walls, whereas the effect is almost negligible for center vials. Variability in heat transfer conditions could result in different crystal structures and drying times, hence variation in the final product quality. Consequently, process control and optimization are not straightforward. These known issues motivate the development of a continuous lyophilization concept that can be practically implemented in an industrial scale.
Continuous lyophilization
A comprehensive review of continuous lyophilization technologies is given in Ref. [5]. To summarize, the current continuous lyophilization technologies can be divided into four main categories. The first category employs equipment that continuously moves a bulk product through the process, e.g., a conveyor [29] or a revolving plate with holes [30]. Another type of technology relies on the concept of spray freeze drying, in which a liquid product is sprayed through nozzles to create small liquid droplets with high surface-to-volume ratio, greatly improving heat and mass transfer in the process [7, 28]. Both aforementioned approaches are proposed for lyophilization of bulk material. For lyophilization of unit doses, a well-known technique uses spin freezing to accelerate the freezing process and improve the uniformity of heat transfer in a product [4].
Finally, the current state-of-the-art continuous lyophilization technology for unit doses has been recently proposed by Ref. [6]. This technology employs the suspended-vial configuration as illustrated in Fig. 1B, where all vials are suspended and continuously moved through the freezing and drying chambers, without any contact between the vials and shelf. This suspended-vial technology has several benefits over the existing ones. First, heat transfer in the equipment is uniform because every vial experiences the exact same heat transfer conditions along the way. In addition, there is a dedicated chamber for controlling the ice nucleation process to help reduce variation in the crystal structure caused by the stochastic nature of ice nucleation. Second, the process does not involve with complicated motions (e.g., high speed flow in spray freeze drying), resulting in a simpler design and more reliable quality control. Third, there is no contact between the vials and cooling/heating shelf, minimizing the risk of having fine particles that could lead to contamination, which is crucial in pharmaceutical applications. Finally, with the automatic filling and load-lock systems, the process is fully continuous. Therefore, our mechanistic model developed in this work is based on this suspended-vial configuration. We refer to Ref. [6] for detailed information and discussion related to the equipment design.
Mechanistic Modeling
This section develops a mechanistic model for continuous lyophilization based on the suspended-vial configuration (see Fig. 1B for a simple schematic and Ref. [6] for detailed discussion). The section starts by discussing the existing models in the literature and then derives the models for all three steps of lyophilization, namely (1) freezing, (2) primary drying, and (3) secondary drying.
Review of the existing models
Mechanistic modeling for conventional batch lyophilization has been studied for many decades, with various models that have been proposed and used in both academia and industry. Despite the fact that the models are for batch processes, modeling strategies underpinning those models can be used as a basis for the development of a continuous lyophilization model for suspended vials. Although there are some previous studies briefly discussing the literature on mechanistic modeling of lyophilization [18, 21], no article systematically summarizes the available modeling strategies and comprehensively discusses pros and cons of each model variation. Hence, this section firstly gives a systematic overview of the mechanistic modeling strategies for lyophilization, discusses pros and cons of each model variation in detail, and finally concludes the optimal model strategies to be used. Note that published models for continuous lyophilization mentioned in Section 1 are based on the spin freezing and spray freeze-drying concepts, which are completely different from the suspended-vial configuration, and so those modeling strategies are not discussed here.
Figure 2: Modeling strategies for the freezing, primary drying, and secondary drying steps in lyophilization. The strategies used in this article are highlighted.
Figure 2 summarizes the modeling strategies for the three different steps in lyophilization, including freezing, primary drying, and secondary drying. Published models mostly focus on primary drying because it is recognized as the most time-consuming, hazardous, and expensive step, and thus the first target for optimization and improvement. A variety of modeling strategies are available, ranging from the simplest lumped capacity model to the high-fidelity multidimensional model. Models for secondary drying have also been well developed as the final product quality (i.e., moisture content) is governed by this step. The mechanistic modeling of freezing receives the least attention in the lyophilization literature due to its complicated behavior, including subcooling and stochastic ice nucleation. As a result, the number of available models for freezing is more limited compared to that of the drying models.
Modeling the freezing step is most complicated among the three steps of localization; the blue column in Fig. 2 highlights modeling strategies for freezing. Due to the limited number of freezing models, we expand our literature search beyond lyophilization to get more complete insights. A conventional technique for modeling the freezing process of pure substance is to apply the concept of a moving boundary problem (aka Stefan problem) where the temperature at the freezing front (solid-liquid interface) is assumed to be constant at the freezing point [31]. Nevertheless, this strategy has not been commonly used in the lyophilization community because it is widely known that the actual freezing process entails subcooling and stochastic ice nucleation, and so the temperature of a product usually goes below the freezing point. A more common but complicated strategy is to consider the subcooling period and incorporate the heat transfer associated with ice nucleation as done in Ref. [8]. Recently, Ref. [11] introduced a state-of-the-art freezing model which incorporates the stochastic nature of ice nucleation [32] and considers heat transfer among multiple vials in batch lyophilization. The same authors subsequently proposed a model for two-dimensional freezing when thermal gradients are significant, e.g., freezing of a product in a large vessel [12]. The most complicated part of freezing is to predict the crystal size or pore size distribution, which directly affects the mass transfer resistance during sublimation in primary drying. While some recent studies have built mechanistic models with the goal of predicting pore size distribution [10, 32], how to best model the specific complex interacting phenomena remains an open research question.
Many models for primary drying are available in the literature (see the green column in Fig. 2). The simplest strategy is to use a lumped capacity model [22], which does not capture the effects of spatial gradients in the system. In both batch and suspended-vial-based continuous lyophilization, vials are heated from the bottom shelf, and the temperature gradient in the vertical direction is substantial. A more common strategy is to model the drying process in one dimension (1D, vertical direction) assuming that the process is controlled by heat transfer only [33, 34, 8, 35]. By omitting mass transfer, the resulting model is simple and has only a few parameters, allowing the analytical solutions to be derived [35], and so the model can be implemented easily. However, as sublimation is a simultaneous heat and mass transfer process, the aforementioned models become inaccurate when the process is controlled by mass transfer. A clever way of incorporating mass transfer into the model while avoiding any unnecessary complexity is to rely on the fact that sublimation occurs at the sublimation front, and so mass transfer can be included as a boundary condition instead of writing a full continuity equation (e.g., see the simplified model of Ref. [18]). Of course, there are a number of models considering detailed heat and mass transfer, both in 1D [14, 15] and in higher dimensions [16]. Nevertheless, implementing those high-fidelity model for practical applications, e.g., state estimation and control, is not easy due to its high complexity, computational time, and number of parameters [18, 21]. Besides, the accuracy of complex models is not significantly different from 1D or simplified models in most cases [21]. Consequently, these high-fidelity models are not commonly used. Note that it is widely acknowledged that the amount of liquid removed via sublimation during primary drying is much higher than that of desorption, and so the effect of desorption can be omitted [15, 16, 18, 21].
Secondary drying is similar to primary drying, with the liquid mainly removed via desorption instead of sublimation. Several modeling strategies are available, including a lumped capacity model [1], 1D heat transfer [25], 1D heat and mass transfer [15], and multidimensional models [16]. In the study by Ref. [25], it was shown that the accuracy of 1D modeling is comparable to that of multidimensional modeling while being much less complicated, whereas a lumped capacity model could produce significant errors. This observation is similar to the primary drying case, where a vertical thermal gradient is usually dominant.
Here we select modeling strategies for our continuous lyophilization model that can capture critical phenomena in the process and produce sufficiently accurate results while keeping the model complexity, computational cost, and number of parameters at a minimum. From the above discussion, the selected strategies are (1) freezing with nucleation and subcooling for the freezing step, (2) 1D heat transfer with mass transfer incorporated as a boundary condition for the primary drying step, and (3) 1D heat transfer with bound water desorption for the secondary drying step, which are also highlighted in Fig. 2. The models for continuous lyophilization developed in the subsequent sections are based on these strategies, with proper modifications for the suspended-vial configuration. Sections 3.2, 3.3, and 3.4 describe the models for freezing, primary drying, and secondary drying, respectively. Additionally, to supplement the mechanistic understanding of the proposed models, Section 3.5 specifically discusses modeling strategies and theories underpinning convection and thermal radiation, the key heat transfer mechanisms in suspended vials.
Figure 3: Schematic diagram showing the mechanistic modeling of suspended vials for (A) freezing, (B) primary drying, and (C) secondary drying.
Model for freezing
For the design proposed by Ref. [6], suspended vials are cooled by using the cryogenic gas at a controlled temperature and flow rate. Besides, a dedicated chamber is added to control the nucleation temperature using the vacuum-induced surface freezing (VISF) technology [36]. With this setup, all surfaces of a vial experience a similar heat transfer condition, which is different from batch lyophilization where only the bottom surface is cooled by the cooling shelf. Given that the vial size is not large and heat transfer at all surfaces is almost identical, it is reasonable to use a lumped capacity model for the freezing step [11].
Our model (Fig. 3A) for freezing of suspended vials is divided into five main steps: (1) preconditioning, (2) VISF, (3) nucleation, (4) solidification, and (5) final cooling.
3.2.1 Preconditioning
In our model, the preconditioning step starts at $t_{0}$ and completes at $t_{\textrm{f1}}$. Preconditioning entails cooling the product in a vial such that the product temperature is uniform at the target value below the freezing point (subcooling). In terms of modeling, this is the simplest step because there is no phase change and nucleation, and so only sensible heat is important.
The energy balance equation for the liquid solution in a vial is
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}},$ (1)
where $T(t)$ is the temperature, $t$ is time, $m$ is the total mass, $C_{p}$ is the specific heat capacity (per mass), $Q$ is the total heat transfer between the vial surface and environment, the subscripts ‘w’ and ‘s’ refer to the water (solvent) and solid (solute) phases, and the subscripts ‘s1’, ‘s2’ and ‘s3’ denote the top, bottom, and side of the vial, respectively. The initial conditions for (1) are
$T(t_{0})=T_{0},$ (2)
where $T_{0}$ is the initial product temperature. During the freezing step in general, the amount of water changes with time due to phase transition, and so we denote $m_{\textrm{w0}}$ as the initial mass of water at $t_{0}$. On the other hand, the amount of solid does not change, meaning that $m_{\textrm{s}}$ is a constant. However, during the preconditioning step, there is no phase transition at all, so
$m_{\textrm{w}}(t_{0}\leq t\leq t_{\textrm{f1}})=m_{\textrm{w,0}}.$ (3)
In some cases, experimental data are reported as the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$, where the subscript ‘l’ denotes the liquid phase (solvent + solute). In such cases, $m_{\textrm{s}}$ and $m_{\textrm{w,0}}$ can be calculated from $V_{\textrm{l}}$ and the densities $\rho_{\textrm{s}}$, $\rho_{\textrm{w}}$ (see Appendix A for the relevant equations).
The next step is to define the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$, which could vary among systems. For example, based on the data reported in Ref. [6], the bottom shelf is used to cool the gas inside the chamber, which subsequently cools the vials, hence natural convection. Besides, there exists thermal radiation between the vials and chamber walls. Therefore, both natural convection and radiation effects are modeled in this case. Consequently, the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are
$Q_{\textrm{s1}}=h_{\textrm{s1}}A_{z}(T_{\textrm{c}}-T),$ (4)
$Q_{\textrm{s2}}=h_{\textrm{s2}}A_{z}(T_{\textrm{g}}-T),$ (5)
$Q_{\textrm{s3}}=h_{\textrm{s3}}A_{z}(T_{\textrm{g}}-T)+\sigma A_{r}\mathcal{F}(T_{\textrm{c}}^{4}-T^{4}),$ (6)
where $h$ is the heat transfer coefficient, $A_{z}$ is the cross sectional area of the product, $A_{r}$ is the side surface area of the product, $\sigma$ is the Stefan-Boltzmann constant, and $\mathcal{F}$ is the transfer factor. In this work, the vial/product is modeled as a cylinder of diameter $d$, and hence the cross sectional and surface area can be calculated from the given volume and diameter. Here there are two important temperatures: chamber wall temperature $T_{\textrm{c}}$ and cold gas temperature $T_{\textrm{g}}$. The top surface (chamber wall) and local gas in that area are not cooled, so we assume the gas temperature is equal to the wall temperature $T_{\textrm{c}}$, with $h_{\textrm{s1}}$ combining the effects of both thermal radiation and natural convection. Similarly, the gas at the bottom surface is cooled by the bottom shelf, so we assume that both gas and bottom shelf have the same temperature $T_{\textrm{g}}$, with $h_{\textrm{s2}}$ combining both radiation and convection. At the side surfaces, there are natural convection from the cold gas at $T_{\textrm{g}}$ and thermal radiation from the chamber wall $T_{\textrm{c}}$, and thus $h_{\textrm{s3}}$ takes into account of the convection part only.
As mentioned before, the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are dependent on the system. Another design proposed by Ref. [6] is to flush the cryogenic gas directly into the chamber. In that case, the dominant heat transfer mode is forced convection, which is much stronger than thermal radiation in this low temperature region. Thus, the effect of thermal radiation might be omitted, simplifying (4)–(6).
3.2.2 Vacuum-induced surface freezing
At the end of the preconditioning step, VISF is initiated at $t_{\textrm{f1}}$ and proceeds until VISF completes at $t_{\textrm{f2}}$. The key idea of VISF is to reduce the total pressure of a system to evaporate the liquid solution, abruptly decreasing the temperature and promoting nucleation. A model for VISF can be derived from the mechanistic understanding of simultaneous heat and mass transfer during evaporation, e.g., as described in Ref. [31]. First, consider the mass transfer part. The evaporation rate of water at the surface for nonvolatile solute is
$\frac{dm_{\textrm{w}}}{dt}=h_{m}A_{z}(x_{\textrm{w,sat}}-x_{\textrm{w,c}}),$ (7)
where $h_{m}$ is the mass transfer coefficient, $x_{\textrm{w,sat}}$ is the mole fraction of water at the liquid-vapor interface (equilibrium), and $x_{\textrm{w,c}}$ is the mole fraction of water in the chamber (environment). The subscripts ‘sat’ and ‘c’ here denote the equilibrium condition and environment, respectively. To calculate the mass fraction of water for (7), we first assume that there are two gas/vapor components during VISF, namely (1) water and (2) nitrogen or inert gas, denoted by the subscript ‘in’. By assuming the ideal gas law,
$x_{\textrm{w,sat}}=\frac{p_{\textrm{w,sat}}M_{\textrm{w}}}{p_{\textrm{w,sat}}M_{\textrm{w}}+(p_{\textrm{t}}-p_{\textrm{w,sat}})M_{\textrm{in}}},$ (8)
$x_{\textrm{w,c}}=\frac{p_{\textrm{w,c}}M_{\textrm{w}}}{p_{\textrm{w,c}}M_{\textrm{w}}+(p_{\textrm{t}}-p_{\textrm{w,c}})M_{\textrm{in}}},$ (9)
where $p_{\textrm{t}}$ is the total pressure, $p_{\textrm{w}}$ is the partial pressure of water, and $M$ is the molar mass. If the environment contains only nitrogen or inert gas, $p_{\textrm{w,c}}=0$. The saturation pressure $p_{\textrm{w,sat}}$ is a function of temperature, that is [37],
$p_{\textrm{w,sat}}=10^{3}\exp\!\left(\!16.3872-\frac{3885.7}{T-42.98}\right).$ (10)
The initial condition for (7) is simply the final temperature of the preconditioning step.
Next, consider the heat transfer part. The energy balance equation can be modified from (1) as
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}(t)C_{p,\textrm{w}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}}(t)+\Delta H_{\textrm{vap}}\frac{dm_{\textrm{w}}}{dt},$ (11)
where the last term on the right-hand side is the amount of heat removed via evaporation and $\Delta H_{\textrm{vap}}$ is the heat of vaporization. At low temperature, $\Delta H_{\textrm{vap}}$ can be approximated by
$\Delta H_{\textrm{vap}}=2.257\mbox{$\times$}10^{6}\!\left(\frac{1-T/647.1}{1-373.15/647.1}\right)^{\!\!0.38}.$ (12)
In (11), $m_{\textrm{w}}(t)$ and $Q_{\textrm{s3}}(t)$ are time-dependent because the amount of water (and hence the side surface area $A_{r}$) changes with time. The initial condition for (11) is the final mass of water at the end of the preconditioning step, which is $m_{\textrm{w,0}}$.
Note that, if the VISF process is done properly, the amount of water vaporized is usually very small and so can be safely neglected.
3.2.3 Nucleation
The nucleation step starts at the end of VISF ($t_{\textrm{f2}}$) for controlled nucleation or at the end of the preconditioning step ($t_{\textrm{f1}}$) for uncontrolled nucleation, and then completes at $t_{\textrm{f3}}$. Our strategy to treat the nucleation step described below is based on the state-of-the-art freezing model proposed by Ref. [11]. In this work, we use the term nucleation to denote the first nucleation where the temperature of a subcooled liquid almost instantaneously increases to the equilibrium/freezing point, whereas the term solidification is used to denote the phase transition from liquid to solid after that first nucleation.
Consider the controlled nucleation case first. The energy balance during nucleation, assuming the process is instantaneous and adiabatic, is
$(T_{\textrm{f,l}}-T_{\textrm{n}})(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}})=m_{\textrm{i,n}}\Delta H_{\textrm{fus}},$ (13)
where $T_{\textrm{f,l}}$ is the freezing point of the liquid solution, i.e., the product temperature after nucleation, $T_{\textrm{n}}$ is the nucleation temperature, i.e., the product temperature when nucleation starts, $m_{\textrm{i,n}}$ is the mass of ice formed immediately after nucleation, the subscript ‘i’ denotes the solid phase (ice), and $\Delta H_{\textrm{fus}}$ is the heat of fusion. With the presence of a non-volatile solute, the freezing-point depression is
$T_{\textrm{f,w}}-T_{\textrm{f,l}}=\frac{K_{\textrm{f}}}{M_{\textrm{s}}}\!\left(\frac{m_{\textrm{s}}}{m_{\textrm{w}}-m_{\textrm{i,n}}}\right),$ (14)
where $T_{\textrm{f,l}}$ is the freezing point of pure water, $K_{\textrm{f}}$ is the molal freezing-point depression constant, and $M_{\textrm{s}}$ is the molar mass of a solute. The two unknowns $T_{\textrm{f,l}}$ and $m_{\textrm{i,n}}$ can be obtained by solving both (13) and (14) simultaneously, both of which become the initials conditions for the next (solidification) step.
Finally, define $t_{\textrm{f3}}$ as the time when the nucleation process completes. Since the nucleation process is nearly instantaneous as explained by Ref. [11], $t_{\textrm{f3}}$ is set to $t_{\textrm{f2}}$, with the relations
$m_{\textrm{w}}(t_{\textrm{f3}})=m_{\textrm{w}}(t_{\textrm{f2}})-m_{\textrm{i,n}},$ (15)
$m_{\textrm{i}}(t_{\textrm{f2}})=0,$ (16)
$m_{\textrm{i}}(t_{\textrm{f3}})=m_{\textrm{i,n}},$ (17)
$T(t_{\textrm{f2}})=T_{\textrm{n}},$ (18)
$T(t_{\textrm{f3}})=T_{\textrm{f,l}}.$ (19)
For the uncontrolled nucleation case (no VISF), the entire calculation is Section 3.2.2 is skipped. By nature, ice nucleation is stochastic and interpreted as a Poisson process. The rate constant of the Poison process $\lambda$ is [32]
$\lambda=k_{\textrm{n}}(T_{\textrm{f,l}}-T)^{b_{\textrm{n}}}V_{\textrm{l}},$ (20)
where $k_{\textrm{n}}$ and $b_{\textrm{n}}$ are the nucleation kinetics parameters. Before nucleation occurs, there is no ice in the system, and hence $T_{\textrm{f,l}}$ can be calculated by
$T_{\textrm{f,w}}-T_{\textrm{f,l}}=\frac{K_{\textrm{f}}m_{\textrm{s}}}{M_{\textrm{s}}m_{\textrm{w}}}.$ (21)
The probability that the first nucleus is formed between time $t$ and $t+\Delta t$ is expressed by
$P=1-\exp(\lambda\Delta t).$ (22)
The probability defined by (22) is calculated cumulatively at every time step $\Delta t$ until the first nucleation occurs, which is marked as the end of the preconditioning step $t_{\textrm{f1}}$. At this point, the exact same procedure described for the controlled nucleation case, i.e.,(13)–(19), is followed, in which $t_{\textrm{f2}}$ is replaced with $t_{\textrm{f1}}$.
3.2.4 Solidification
After the first nucleation completes, the solidification step starts at $t_{\textrm{f3}}$. During this step, the product temperature is equal to the freezing point of the solution, which follows the freezing-point depression relation
$T(t)=T_{\textrm{f,w}}-\frac{K_{\textrm{f}}}{M_{\textrm{s}}}\!\left(\frac{m_{\textrm{s}}}{m_{\textrm{w}}(t)}\right),$ (23)
The energy balance during the freezing step is
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}(t)C_{p,\textrm{w}}+m_{\textrm{i}}(t)C_{p,\textrm{i}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}(t)+Q_{\textrm{s3}}+\Delta H_{\textrm{fus}}\frac{dm_{\textrm{i}}}{dt},$ (24)
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (4), (5), and (6), respectively. During the solidification step, the amount of water and ice changes with time, which affects the volume and surface area of the product, and so $Q_{\textrm{s2}}(t)$ is time-dependent. The relation between the amount of water and ice is
$m_{\textrm{w}}(t)=m_{\textrm{w}}(t_{\textrm{f3}})-m_{\textrm{i}}.$ (25)
The initial conditions for $m_{\textrm{w}}$, $m_{\textrm{i}}$, and $T$ are the final conditions of the nucleation step, namely (15), (17), and (19).
Finally, define $t_{\textrm{f4}}$ as the time when solidification completes. The criterion for complete solidification is defined as
$m_{\textrm{i}}(t_{\textrm{f4}})=0.95m_{\textrm{w}}(t_{\textrm{f3}}),$ (26)
where the coefficient could vary between 0.85 and 0.95 without significantly changing the final results [11].
3.2.5 Cooling
The final step is to ensure that the product temperature is at the desire value before starting the drying process. This cooling step starts at $t_{\textrm{f4}}$ and completes at $t_{\textrm{f5}}$. The model for the final cooling step is relatively simple, with the only energy balance equation
$(m_{\textrm{s}}C_{p,\textrm{s}}+m_{\textrm{w}}C_{p,\textrm{w}}+m_{\textrm{i}}C_{p,\textrm{i}})\frac{dT}{dt}=Q_{\textrm{s1}}+Q_{\textrm{s2}}+Q_{\textrm{s3}},$ (27)
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (4), (5), and (6), respectively. Since there is no phase change during this step, the amount of substances, volume, and surface area are all constant, taking the final conditions of the solidification step. The initial condition for (27) is the final temperature of the solidification step. It is generally known that, after solidification, the amount of bound water is negligible compared to that of ice, and so $m_{\textrm{w}}C_{p,\textrm{w}}$ could be omitted from (27) without any significant error.
By consecutively simulating the models developed in Sections 3.2.1 to 3.2.5, the evolution of the product temperature and amount of substances (e.g., water and ice) during the freezing
Model for primary drying
In conventional lyophilization where vials are placed on the heating shelf, there are three heat transfer mechanisms associated with the bottom surface, namely conduction, convection, and radiation, whereas only radiation dominates at the side and top surfaces [17]. For continuous lyophilization, suspended vials are heated by the below heating shelf without any contact between the vials and shelf. As a result, heat transfer at the bottom surface is driven only by thermal radiation and natural convection.
The model for primary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) (Fig. 3B). Define the primary drying step to start at $t_{0}$ and complete at $t_{\textrm{d1}}$. If the primary drying model is simulated following the freezing step, then set $t_{0}=t_{\textrm{f4}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone primary drying simulation.
The governing equations for the primary drying step consists of (1) the energy balance in the frozen region and (2) the mass balance at the sublimation front. By assuming that the supplied heat is used in the frozen region only, the energy balance for the frozen region is
$\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{\partial T}{\partial t}=k_{\textrm{f}}\dfrac{\partial^{2}T}{\partial z^{2}}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad S(t)<z<H,$ (28)
where $T(z,t)$ is the temperature, $S(t)$ is the sublimation front/interface position, $k$ is the thermal conductivity, $\rho$ is the density, $C_{p}$ is the heat capacity, $V$ is the volume, $H$ is the height of the product, and the subscript ‘f’ denotes the frozen region. We refer to Appendix A for the calculations of some relevant parameters. The radiative heat transfer from the sidewall $Q_{\textrm{rad}}$ is
$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}\!\left(T_{\textrm{c}}^{4}-T^{4}\right),$ (29)
where $\mathcal{F}$ is the transfer factor, $\sigma$ is the Stefan-Boltzmann constant, $T_{\textrm{c}}$ is the chamber wall temperature, and $A_{r}=\pi dH$ is the side area of the product. We refer to Section 3.5 for the detailed derivation of (29). The mass balance of water at the sublimation front gives
$\frac{dS}{dt}=\frac{N_{\textrm{w}}}{\rho_{\textrm{f}}-\rho_{\textrm{e}}},$ (30)
where $N_{\textrm{w}}$ is the sublimation flux and $\rho_{\textrm{e}}$ is the effective density of the dried region above the sublimation front. The driving force for mass transfer at the sublimation interface is [17, 1, 22]
$N_{\textrm{w}}=\frac{p_{\textrm{w,sat}}-p_{\textrm{w,c}}}{R_{\textrm{p}}},$ (31)
where $p_{\textrm{w,sat}}$ is the saturation/equilibrium pressure of water, $p_{\textrm{w,c}}$ is the partial pressure of water in the chamber (environment), and $R_{\textrm{p}}$ is the mass transfer resistance. The saturation pressure for sublimation can be is described by [22]
$p_{\textrm{w,sat}}=\exp\!\left(\frac{-6139.9}{T}+28.8912\right).$ (32)
The variation of the resistance normally follows the empirical expression [17, 21, 22]
$R_{\textrm{p}}=R_{\textrm{p0}}+\frac{R_{\textrm{p1}}S}{R_{\textrm{p2}}+S},$ (33)
where $R_{\textrm{p}}$ is the mass transfer resistance (aka cake resistance) and $R_{\textrm{p0}}$, $R_{\textrm{p1}}$, $R_{\textrm{p2}}$ are the constants estimated from data.
The partial differential equation (PDE) (28) requires two boundary conditions. Heat transfer at the bottom surface of the frozen product follows Newton’s law of cooling,
$-k_{\textrm{f}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}(t)),\qquad z=H,$ (34)
where $T_{\textrm{b}}$ is the bottom shelf temperature and $h_{\textrm{b}}$ is the overall heat transfer coefficient which combines the effects of thermal radiation and natural convection. At the top surface, the energy balance associated with sublimation and thermal radiation from the above shelf/wall is
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=k_{\textrm{f}}\dfrac{\partial T}{\partial z}+\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad z=S(t),$ (35)
where $\Delta H_{\textrm{sub}}$ is the heat of sublimation and $\varepsilon_{\textrm{gl}}$ is the emissivity of the glass vial.
The initial conditions for (28) and (30) are
$T(z,t_{0})=T_{0},\qquad 0\leq z\leq H,$ (36)
$S(t_{0})=0.$ (37)
If the model is simulated following the freezing step, then set $T_{0}=T(t_{\textrm{f4}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone primary drying simulation.
The primary drying model is simulated until the interface position is equal to the height of the product $S(t)=H$, indicating that there is no frozen material left, which marks the end of the primary drying step $t_{\textrm{d1}}$. In some cases, there might be an additional heating period at the end of primary drying to adjust the temperature and ensure complete sublimation before starting the secondary drying stage. However, this period is trivial to model; only a simple heat equation is required, and so is not described it here.
Model for secondary drying
The model for secondary drying is formulated in the rectangular coordinate system with one spatial dimension ($z$) and time ($t$) as done for the primary drying model (Fig. 3C). The secondary drying step is defined to start at $t_{0}$ and complete at $t_{\textrm{d2}}$. If the secondary drying model is simulated following the primary drying step, then set $t_{0}=t_{\textrm{d2}}$. Otherwise, $t_{0}$ should be set to 0 for a standalone secondary drying simulation.
The governing equations for the secondary drying step consists of (1) the energy balance in the dried region and (2) the desorption kinetics. The energy balance of the dried product is
$\rho_{\textrm{e}}C_{p,\textrm{e}}\dfrac{\partial T}{\partial t}=k_{\textrm{e}}\dfrac{\partial^{2}T}{\partial z^{2}}+\rho_{\textrm{d}}\Delta H_{\textrm{des}}\dfrac{\partial c_{\textrm{w}}}{\partial t}+\frac{Q_{\textrm{rad}}}{V_{\textrm{e}}},\qquad 0\leq z\leq H,$ (38)
where $T(z,t)$ is the product temperature, $c_{\textrm{w}}(z,t)$ is the bound water concentration (aka moisture content, residual water), $\rho_{\textrm{d}}$ is the density of the dried region (solid and vacuum), $\Delta H_{\textrm{des}}$ is the enthalpy of desorption, $Q_{\textrm{rad}}$ is as defined in (29), and the subscript ‘e’ denotes the effective properties considering both solid and gas in the pores. The desorption kinetics of bound water is described by
$\dfrac{\partial c_{\textrm{w}}}{\partial t}=k_{\textrm{d}}(c^{*}_{\textrm{w}}-c_{\textrm{w}}),$ (39)
where $c^{*}_{\textrm{w}}$ is the equilibrium concentration of bound water and $k_{\textrm{w}}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [14, 15, 21]
$k_{\textrm{d}}=f_{\textrm{a}}e^{-E_{\textrm{a}}/RT},$ (40)
where $f_{\textrm{a}}$ is the frequency factor (aka collision frequency) and $E_{\textrm{a}}$ is the activation energy. The above desorption kinetics (39) is known as the linear driving force model, one of the simplest adsorption/desorption models that can accurately predict the dynamics of bound water and has been widely used in the literature [14, 38, 18, 23, 21]. Further simplification that is very common in the literature can be done by setting $c^{*}_{\textrm{w}}=0$. This simplification produces insignificant error as shown in Ref. [15] and eliminates the need for equilibrium data and detailed knowledge about the solid structure [21].
The governing PDE (38) requires two boundary conditions. The bottom surface of the dried product is heated by the heating shelf, which follows Newton’s law of cooling
$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=h_{\textrm{b}}(T-T_{\textrm{b}}(t)),\qquad z=H.$ (41)
Heat transfer at the top surface is mainly thermal radiation, resulting in the boundary condition
$-k_{\textrm{e}}\dfrac{\partial T}{\partial z}=\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad z=0.$ (42)
The initial conditions for (38) and (39) are
$T(z,t_{0})=T_{0},\qquad 0\leq z\leq H,$ (43)
$c_{\textrm{w}}(z,t_{0})=c_{\textrm{w,0}},\qquad 0\leq z\leq H.$ (44)
If the model is simulated following the primary drying step, then set $T_{0}=T(z,t_{\textrm{d1}})$. Otherwise, $T_{0}$ can be set arbitrarily for a standalone secondary drying simulation.
The secondary drying model should be simulated until the concentration of bound water is below the target value, denoted as $c_{\textrm{w,}\infty}$, which marks the end of the secondary drying step $t_{\textrm{d2}}$.
Modeling heat transfer in suspended vials
Convection and radiation are two key heat transfer mechanisms in the suspended-vial configuration as discussed before. During the freezing step, there could be a combination of natural/forced convection and radiation, depending on how the cryogenic gas is fed to the system. In primary and secondary drying, only natural convection and radiation are important.
Convection can be simply modeled using Newton’s law of cooling. A heat transfer coefficient for convection can be estimated from correlations that entail dimensionless groups such as the Nusselt number, Prandtl number, Reynolds number (forced convection), and Grasholf number (natural convection). Nevertheless, it is more common and accurate to estimate the heat transfer coefficient from data, e.g., using the techniques suggested in Ref. [21].
Thermal radiation is significantly more complicated in terms of modeling. The theories and model equations described below are mainly based on Refs. [31] and [39]; the former discusses general thermal radiation theories, while the latter provides a detailed study on the modeling of thermal radiation in lyophilization. Here we summarize only the key elements needed for modeling thermal radiation in the suspended-vial configuration; please refer to the aforementioned references for more details. Consider radiation exchange between two diffuse, gray surfaces of finite size denoted as surfaces 1 and 2, respectively. In general, the net radiant energy receiving by surface 1 can be written as
$Q_{\textrm{rad}}=\sigma A_{1}\mathcal{F}(T_{2}^{4}-T_{1}^{4}),$ (45)
where $A$ is the surface area, $\varepsilon$ is the emissivity, $\mathcal{F}$ is the transfer factor, and the subscripts ‘1’ and ‘2’ denote surfaces 1 and 2, respectively. The transfer factor is dependent on geometry and material type (emissivity). For any two diffuse, gray surfaces that form an enclosure,
$Q_{\textrm{rad}}=\dfrac{\sigma(T_{2}^{4}-T_{1}^{4})}{\dfrac{1-\varepsilon_{1}}{\varepsilon_{1}A_{1}}+\dfrac{1}{A_{1}F_{1-2}}+\dfrac{1-\varepsilon_{2}}{\varepsilon_{2}A_{2}}},$ (46)
where $F_{1-2}$ is the view/shape factor. If surface 1 is surrounded by surface 2 ($F_{1-2}=1$) and surface 2 is much larger than surface 1, (46) can be simplified as
$Q_{\textrm{rad}}=\varepsilon_{1}\sigma A_{1}(T^{4}_{2}-T^{4}_{1}).$ (47)
In (47), the transfer factor $\mathcal{F}=\varepsilon_{1}$. Now consider the case when surface 1 is the vial and surface 2 is the chamber walls. In this case, (47) becomes
$Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{r}(T^{4}_{\textrm{c}}-T^{4})$ (48)
for the side surface and
$Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{z}(T^{4}_{\textrm{c}}-T^{4})$ (49)
for the top and bottom surfaces. Here the transfer factor $\mathcal{F}=\varepsilon_{\textrm{gl}}$. The amount of radiative heat for the single-vial case represented by (48) and (49) defines the upper bound on the radiative heat that the vial can receive from the chamber walls. This is because, when there are multiple vials, the view factor $F_{1-2}$ is less than 1, and so the radiative heat is shared among vials. As a result, the transfer factor $\mathcal{F}$ for the multiple-vial case must be less than $\varepsilon_{\textrm{gl}}$. Therefore, (48) needs to be modified to account for interactions between vials, whereas (49) can be used directly because thermal radiation at the top and bottom surfaces is not influenced by the other vials.
Typical lyophilization of unit doses always consists of a large number of vials, and so thermal radiation exchange exists not only between vials and chamber walls but also between those multiple vials. In batch lyophilization, vials are arranged as an array consisting of many rows, which results in significant differences in heat transfer conditions between the outer and inner vials. In such cases, the outer vials with higher temperature transfer significant heat to the inner vials while simultaneously exchanging heat with the chamber walls. A rigorous way of modeling this complicated radiation is known as the radiation network approach, which is comprehensively discussed and demonstrated in the context of lyophilization in Ref. [39]. In the suspended-vial configuration, however, vials are typically aligned in a few rows, e.g., a single row in the equipment proposed by Ref. [6]. As a result, all vials experience nearly the same heat transfer heat condition, resulting in negligible heat transfer between vials. Therefore, instead of using the radiation network approach, we modify (48) to follow (45) as
$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}(T^{4}_{\textrm{c}}-T^{4}),$ (50)
where $\mathcal{F}$ must be less than $\varepsilon_{\textrm{gl}}$ as described in the previous paragraph. This transfer factor $\mathcal{F}$ is best to be estimated from experimental data but can also be approximated mechanistically, e.g., using (46). This technique is similar to the hybrid approach presented in Ref. [39]. Note that this approximation is for the suspended-vial configuration, in which heat transfer between vials is not significant; otherwise, the radiation network approach is recommended.
When there exists both convection and radiation, it is more convenient to linearize and rewrite the radiation part in the form of Newton’s law of cooling; for the temperature of about 300K, the error is about 0.1% [31]. The linearized equation can then be combined with the convection part, with the overall heat transfer coefficient taking into account of both convection and radiation, e.g., (34).
[…]
Default model parameters
The proposed model developed in Sections 3.2, 3.3, and 3.4 require a large number of input parameters. This section defines the default parameter values for our model to be used in the subsequent sections in this work. These parameter values are either obtained from the literature or set based on the typical values used in lyophilization. This default set of parameters assumes a complete simulation in which the models for freezing (with controlled nucleation), primary drying, and secondary drying are simulated consecutively. Table LABEL:Tab:DefaultParameters shows the default model parameters; parameter values different from those reported in the table are stated explicitly in each section or case study.
Table 1: Default model parameters for a complete continuous lyophilization simulation.
Symbol Value Unit Source
Thermophysical properties
[…]
$\Delta H_{\textrm{des}}$ 2.68$\times$106 J/kg [15]
$\Delta H_{\textrm{fus}}$ 3.34$\times$105 J/kg [37]
$\Delta H_{\textrm{sub}}$ 2.84$\times$106 J/kg [18]
$\Delta H_{\textrm{vap}}$ see (12) J/kg [37]
[…]
$\rho_{\textrm{e}}$ 215 kg/m3 [15]
$\rho_{\textrm{d}}$ 212.21 kg/m3 [15]
$\rho_{\textrm{f}}$ 937 kg/m3 calculated (see Appendix A)
$\rho_{\textrm{i}}$ 917 kg/m3 [32]
$\rho_{\textrm{s}}$ 1,587.9 kg/m3 [32]
$\rho_{\textrm{w}}$ 1,000 kg/m3 [12]
Operating conditions
$c_{\textrm{w,0}}$ 0.088 kg water/kg solid [6]
$c_{\textrm{w,}\infty}$ 0.01 kg water/kg solid [1]
$p_{\textrm{t}}$ see Fig. 5D Pa
$p_{\textrm{w,c}}$ 3 Pa [6]
$p_{\textrm{w,sat}}$ see (10) and (32) Pa [37, 22]
$T_{0}$ 298.15 K
$T_{\textrm{b}}$ 270 (primary drying) K
295 (secondary drying) K
$T_{\textrm{c}}$ 273 (before VISF) K
240 (after VISF) K
265 (primary drying) K
290 (secondary drying) K
$T_{\textrm{f,w}}$ 273.15 K [12]
$T_{\textrm{g}}$ 268 (before VISF) K
230 (after VISF) K
$T_{\textrm{n}}$ 268 K
Heat and mass transfer
$\mathcal{F}$ 0.78
$f_{\textrm{a}}$ 1.5$\times$10-3 1/s
[…]
$h_{\textrm{m}}$ 6.34$\times$10-3 kg/m2$\cdots$ [31]
[…]
$R_{\textrm{p0}}$ 1.5$\times$104 m/s estimated from data in Ref. [6]
$R_{\textrm{p1}}$ 3.4$\times$107 1/s estimated from data in Ref. [6]
$R_{\textrm{p2}}$ 1 1/m estimated from data in Ref. [6]
[…]
Product and formulation
$H$ 7.2$\times$10-3 m calculated (see Appendix A)
$m_{\textrm{s}}$ 1.53$\times$10-4 kg calculated (see Appendix A)
$m_{\textrm{w,0}}$ 2.9$\times$10-3 kg calculated (see Appendix A)
$V_{\textrm{l}}$ 3$\times$10-6 m3 [6]
$x_{\textrm{s}}$ 0.05 [6]
Vial
$d$ 0.024 m 10R vial
$\epsilon_{\textrm{gl}}$ 0.8 [31]
Others
[…]
$M_{\textrm{in}}$ 28$\times$10-3 kg/mol [40]
$M_{\textrm{s}}$ 0.3423 kg/mol [12]
$M_{\textrm{w}}$ 18$\times$10-3 kg/mol [40]
[…]
Parameters that should be estimated from experimental data include heat transfer coefficients, mass transfer coefficients and cake resistance, and desorption-related parameters (frequency factor and activation energy).
Implementation
In this article, the mechanistic model was implemented in MATLAB R2023a, and so all results and plots were generated using MATLAB. A system of ordinary differential equations (ODEs) was integrated with ode15s. Partial differential equations (PDEs) were discretized spatially to produce a system of ODEs, which was then fed to the ODE solver (see Appendix B for the detailed numerical methods).
Model validation
The mechanistic model presented in Section 3 is validated using the experimental data reported in Ref. [6]. The model is validated for each lyophilization step separately, with the parameters reported in Table 2. The default values defined in Table LABEL:Tab:DefaultParameters are used for parameters not shown in Table 2.
Table 2: Parameters used for the model validation, including Cases 1, 2a, 2b, 3a, and 3b.
Symbol Value Unit Source
Freezing (no VISF)
$T_{0}$ 280 K [6]
$T_{\textrm{c}}$ 272 K estimated from data in Ref. [6]
$T_{\textrm{g}}$ see Fig. 4A K [6]
$T_{\textrm{n}}$ 263.18 K [6]
[…]
Primary drying
$T_{0}$ 231 (Case 2a) K [6]
235 (Case 2b) K [6]
$T_{\textrm{b}}$ 263 (Case 2a) K [6]
313 (Case 2b) K [6]
$T_{\textrm{b}}$ 275 K estimated from data in Ref. [6]
[…]
$R_{\textrm{p0}}$ 1.5$\times$104 m/s estimated from data in Ref. [6]
$R_{\textrm{p1}}$ 3.4$\times$107 1/s estimated from data in Ref. [6]
$R_{\textrm{p2}}$ 1 1/m estimated from data in Ref. [6]
Secondary drying
$c_{\textrm{w,0}}$ 0.088 (Case 3a) kg water/kg solid [6]
0.075 (Case 3b) kg water/kg solid [6]
$T_{0}$ 273 K [6]
$T_{\textrm{b}}$ 293 K [6]
$T_{\textrm{c}}$ 285 K estimated from data in Ref. [6]
$f_{\textrm{a}}$ 0.42 1/s estimated from data in Ref. [6]
[…]
$V_{\textrm{l}}$ 2$\times$10-6 m3 [6]
Figure 4: Model validation for the lyophilization of suspended vials using the experimental data from Ref. [6]. Panel A shows the model prediction and experimental data for the product temperature during the freezing step. Panel B shows the model prediction and experimental data for the product temperature (assumed to be measured at the bottom surface) during the primary drying step. The maximum shelf temperatures for Cases 2a and 2b are 263 and 313 K, respectively. Panel C shows the model prediction and experimental data for the bound water concentration during the secondary drying step. The initial concentrations for Cases 3a and 3b are 0.0864 and 0.0739 kg water/kg solid, respectively.
For the freezing step, the simulated temperature and freezing time agree well with the experimental data (Fig. 4A). During the preconditioning phase, the temperature decreases from its initial value of about 280 K to the nucleation temperature of about 263 K, where nucleation starts. At the first nucleation, the temperature instantaneously increases from 263 K to about the equilibrium freezing point of about 272 K. The temperature slowly decreases during the solidification phase, which takes about 30 min, following the freezing-point depression. Subsequently, the product temperature decreases and becomes equilibrium with the environment (chamber wall and cryogenic gas), with the simulated temperature approaching the equilibrium slightly faster than the actual temperature.
The model can accurately predict the evolution of the product temperature and drying time during primary drying for both low (263 K) and high (313 K) shelf temperatures (Fig. 4B). The product temperature increases fast at the beginning because the driving force for sublimation is low, and so most of the heat input is used to increase the temperature. As time progresses, the driving force for sublimation is higher, resulting in most of the heat used for sublimation instead of increasing the product temperature. The maximum error in temperature prediction is about 2–3 K, while that of drying time prediction is about 0.2 for Case 2a and 0.4 for Case 2b ($\sim$ 6%). This error could be partly because of the uncertainty of cake resistance, which is caused by a stochastic freezing process.
In secondary drying, the simulated concentration of bound water closely aligns with the experimental data for both concentration levels (0.0864 and 0.0739 kg water/kg solid). The concentration decreases exponentially following the linear driving force model, where the simulations are terminated when the concentration reaches 0.01 kg water/kg solid. It is important to note that the data used in Fig. 4C are based on the shelf temperature of 293 K. To properly estimate the values of $f_{\textrm{a}}$ and $E_{\textrm{a}}$, it is recommended to obtain the concentration profiles at various shelf temperatures such that the effects of temperature on the desorption process can be quantified.
Case Studies
The previous sections detail the development and validation of our mechanistic model for lyophilization. This section demonstrates various applications of our model via four different case studies.
Process design and optimization
The first case study focuses on model-based design and optimization of the entire continuous lyophilization process. One of the critical aspects of continuous manufacturing is to ensure that the process is operated smoothly, optimally, and safely without human intervention, to minimize downtime and maximize production. In the context of continuous lyophilization, our model can be used to study the evolution of critical process parameters, e.g., temperature, moisture content, throughout the process at various operating conditions, which can help guide the process design and optimization.
This work not only presents the first mechanistic model for continuous lyophilization, but also is one of the very few studies that proposes a complete model that includes all three steps in lyophilization, including freezing, primary drying, and secondary drying. This is important specifically for continuous manufacturing where the entire process should be designed and optimized simultaneously.
By using the default parameters in Table LABEL:Tab:DefaultParameters, we simulate a complete continuous lyophilization cycle including (1) freezing with controlled nucleation via VISF, (2) primary drying, and (3) secondary drying (Fig. 5). The evolution of the product temperature (Fig. 5A), mass of ice (Fig. 5B), and bound water concentration (Fig. 5C) can be obtained, given the operating pressure (Fig. 5D) and temperature (Fig. 5E). This information could be useful for various purposes. For example, it can guide the process and equipment design concerning chamber sizing and vial moving speed to accommodate the residence time of each step. Simulation results can also be used to help identify pressure and temperature profiles to achieve some specific objectives, e.g., minimization of drying time. It is important to note that the entire simulation shown in Fig. 5 can be computed in less than 1 second on a normal laptop, which makes the model very practical to be implemented for any purposes.
Figure 5: Complete simulation for the suspended-vial continuous lyophilization process showing the time evolution of the (A) product average temperature, (B) total mass of ice, (C) concentration of bound water, (D) operating pressure, and (E) operating temperature.
Studying vacuum-induced surface freezing
Our model can be simulated for the whole lyophilization process as shown in Section 5.1 or simulated separately for each step in the lyophilization cycle to provide insights into some part of the process. This section employs the model to study the freezing step, which usually receives little attention compared to the drying steps.
The proposed model can simulate different freezing protocols, namely uncontrolled (aka spontaneous) and controlled nucleation. Spontaneous nucleation was already investigate in detail by Refs. [11, 12], and so here we consider the controlled nucleation case via VISF, which is preferable for continuous lyophilization. As an example, the evolution of the product temperature and amount of water evaporating during VISF at different pressure profiles is simulated (Fig. 6). Varying among different systems, this information could be used to design a freezing protocol, e.g., pressure profile, that achieves the desired nucleation temperature.
Figure 6: Simulation results showing the evolution of the product during VISF at different pressure profiles. Panel A shows the pressure profiles that decrease linearly from $10^{5}$ Pa to $10^{4}$ Pa, 1,000, and 100 Pa, respectively. Panel B shows the evolution of the product temperature. Panel C shows the total mass of water, which implies the evaporation rate. External heat transfer with the wall and cryogenic gas is excluded so that only the contribution from VISF is considered.
Model modifications for conventional drying
Although the development of the mechanistic model in this article is for continuous lyophilization, the model can be easily modified for batch/conventional lyophilization. The major modifications are to turn off the thermal radiation part in the model and recalculate/estimate the heat transfer coefficient and cake resistance. In this case study, we consider batch primary drying and secondary drying using the data from Refs. [18] and [21], with the parameters reported in Table 3.
Table 3: Parameters used for the conventional drying studies, including for Cases 4a, 4b, and 5.
Symbol Value Unit Source
Primary drying
[…]
$\rho_{\textrm{e}}$ 252 kg/m3 [18]
$\rho_{\textrm{f}}$ 917 kg/m3 [18]
$p_{\textrm{w,c}}$ 15 Pa [18]
$T_{0}$ 228.15 K [18]
$T_{\textrm{b}}$ see1 K [18]
$\mathcal{F}$ 0 turn off the radiation part
[…]
$R_{\textrm{p0}}$ 3.8$\times$104 m/s estimated from data in Ref. [18]
$R_{\textrm{p1}}$ 3.1$\times$107 1/s estimated from data in Ref. [18]
$R_{\textrm{p2}}$ 10 1/m estimated from data in Ref. [18]
$H$ 7.15$\times$10-3 m [18]
Secondary drying
$c_{\textrm{w,0}}$ 0.0603 kg water/kg solid [21]
$T_{0}$ 273 K [21]
$T_{\textrm{b}}$ see2 K estimated from data in Ref. [21]
$f_{\textrm{a}}$ 1.2$\times$10-3 1/s estimated from data in Ref. [21]
[…]
$\mathcal{F}$ 0 turn off the radiation part
[…]
$V_{\textrm{l}}$ 1.5$\times$10-6 m3 [21]
$d$ 0.01425 m 2R vial, [21]
  • 1
    $T_{\textrm{b}}$ increases linearly at the rate of 0.25 K/min from 228.15 K to 258.15 K (Case 4a) and 268.15 (Case 4b).
  • 2
    $T_{\textrm{b}}$ increases linearly at the rate of 0.5 K/min from 264.09 K to 312 K for Case 5.
Our model provides reliable estimates of the product temperature and sublimation interface position during primary drying at two different shelf temperatures (Fig. 7). The model can also effectively predict the product temperature and bound water concentration during the secondary drying step (Fig. 8). With proper modifications, our continuous lyophilization model is capable of simulating conventional lyophilization.
Figure 7: Comparison between our model prediction and the experimental/simulation data from Ref. [18] for the primary drying step in conventional lyophilization. Panel A shows the model prediction and the experimental data for the product temperature at the bottom surface. Panel B shows the model prediction and reference simulation data for the sublimation front position. The maximum shelf temperatures for Cases 4a and 4b are 258.15 and 268.15, while other conditions are identical.
Figure 8: Comparison between our model prediction and the experimental data from Ref. [21] for the secondary drying step in conventional lyophilization. Panel A shows the model prediction and experimental data for the product temperature at the bottom surface. Panel B shows the model prediction and experimental data for the concentration of bound water.
Condenser failure analysis
Typical process modeling focuses on behaviors of the system during its normal operation where the process is operated steadily at its desired conditions. A better process model should be able to simulate important abnormal conditions, e.g., equipment failure, that could occur due to various unexpected scenarios, which is critical in continuous manufacturing as these abnormal operations could affect the reliability, availability, and maintainability (RAM) as well as the safety of the plant/process.
In lyophilization, one of the crucial design considerations is associated with a condenser. During normal operation, the maximum capacity of a condenser must be higher than the rate of vapor production via sublimation, and choked flow should be avoided [1]. This is to ensure that there will not be vapor accumulation in the equipment, which could subsequently lead to in a pressure increase. This scenario may not lead to safety-related issues because the operating pressure in lyophilization is usually low and the amount of ice/water is not high enough to create overpressure. Nevertheless, an increase in the total pressure reduces the driving force for sublimation (see (31)), which then prolongs the primary drying stage. As a result, the final product quality could be affected. This section explores how to incorporate those dynamic behaviors into our model.
Most primary drying models, including our model, assume that the system pressure is well controlled, which is a reasonable assumption. To incorporate the condenser and dynamics of vapor flow into the model, consider the mass balance equation for water vapor in a chamber [22]
$\frac{dp_{\textrm{w,c}}}{dt}=\frac{(j_{\textrm{w}}-j_{\textrm{w,max}})R\overline{T}}{V_{\textrm{c}}M_{\textrm{w}}},$ (51)
where $j_{\textrm{w}}$ is the total mass flow rate of water vapor, $j_{\textrm{w,max}}$ is the maximum condenser capacity, $V_{\textrm{c}}$ is the volume of a chamber, and $T_{\textrm{avg}}$ is the average temperature in the chamber assumed to be constant for simplification. The mass flow rate of water vapor can be calculated by
$j_{m,\textrm{w}}=n_{\textrm{vial}}A_{z}N_{w},$ (52)
where $n_{\textrm{vial}}$ is the number of vials in the chamber. By coupling (51) and (52) with the primary drying model developed in Section 3.3, the effects of condenser failure or choked flow can be quantified.
Here we consider a situation where there is some issue with the condenser, and so the total condenser capacity cannot handle the water vapor produced via sublimation during primary drying, with the parameters given in Table 4. The simulation result shows that there is a pressure increase from the normal operating value of 3 Pa to 20 Pa (Fig. 9A). This pressure increase results from the total vapor flow that exceeds the maximum condenser capacity, leading vapor accumulation in the chamber. When the pressure increases, the driving force for sublimation (also the sublimation flux) decreases. Finally, when the sublimation flux balances with the condenser capacity, the system pressure becomes constant. It is also shown that the drying time increases by about an hour (Fig. 9B), and the product temperature is higher (Fig. 9C).
Table 4: Parameters used for the condenser failure analysis.
Symbol Value Unit Source
$T_{0}$ 228 K
$T_{\textrm{b}}$ 270 K
$\overline{T}$ 260 K assumed to be constant
$n_{\textrm{vial}}$ 200
$V_{\textrm{c}}$ 0.118 m3 [22]
$j_{m,\textrm{max}}$ 1.8$\times$10-5 m3 [22]
Figure 9: Simulation results comparing the normal operation and condenser failure scenario for the (A) system pressure, (B) sublimation front position, and (C) product average temperature during primary drying.
Conclusion
This work proposes the first mechanistic model for the suspended-vial continuous lyophilization technology, in which vials containing the pharmaceutical products are suspended and moved continuously along the process. The model can be used to simulate the entire lyophilization process, including the freezing, primary drying, and secondary drying steps. The freezing model considers preconditioning, vacuum-induced surface freezing (VISF), spontaneous and controlled nucleation, solidification, and final cooling. The primary drying model captures heat transfer in the frozen region and ice sublimation driven by mass transfer at the sublimation front. The secondary drying model describes simultaneous heat transfer in the dried region and bound water desorption. The overall model can describe the evolution of product temperature, ice/water fraction, and residual moisture throughout the process.
Our model is validated using experimental data from the literature, in which the model predictions are consistent with the experimental measurements in all cases. The validated model is applied for four case studies that entail continuous lyophilization process design and optimization, detailed investigation of VISF, model modifications for conventional drying, and condenser failure analysis. The framework and results presented in this work are promising and could be used to guide the design and development of future continuous lyophilization technology.
With a well-developed mechanistic model, future work could consider utilizing the model for several crucial tasks, e.g., uncertainty quantification, state estimation, and control design, which will ultimately lead to the development of a high-quality digital twin for continuous lyophilization that can greatly facilitate the process design and development in this area.
Data Availability
All software and data will be made available upon publication of the manuscript.
Acknowledgements
This research was supported by the U.S. Food and Drug Administration under the FDA BAA-22-00123 program, Award Number 75F40122C00200.
Appendices
Appendix A Supplemental equations
This appendix summarizes additional equations that are not crucial part of the model but required for calculating some model parameters.
In Section 3.2.1, if the total liquid volume $V_{\textrm{l}}$ and mass fraction of a solute $x_{\textrm{s}}$ are reported, the density of the liquid phase $\rho_{\textrm{l}}$ is
$\rho_{\textrm{l}}=\frac{V_{\textrm{l}}}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{w}}}}.$ (A.1)
Consequently, $m_{\textrm{s}}$ and $m_{\textrm{w,0}}$ can be calculated by
$m_{\textrm{s}}=x_{\textrm{s}}\rho_{\textrm{l}}V_{\textrm{l}},$ (A.2)
$m_{\textrm{w,0}}=(1-x_{\textrm{s}})\rho_{\textrm{l}}V_{\textrm{l}}.$ (A.3)
Then, (1) can be used as usual.
During the freezing step, the surface area $A_{r}$ changes with time due to changes in the total volume caused by phase transition. In general, the surface area can be calculated by
$A_{r}(t)=\frac{4\!\left(\dfrac{m_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{m_{\textrm{w}}(t)}{\rho_{\textrm{w}}}+\dfrac{m_{\textrm{i}}(t)}{\rho_{\textrm{i}}}\right)}{d},$ (A.4)
where the product is assumed to be a cylinder.
For the primary drying model (Section 3.3), the volume of the frozen region is
$V_{\textrm{f}}=\pi d^{2}(H-S).$ (A.5)
The properties of the frozen region (subscript ‘f’) can be calculated from those of the ice and solid material or solute as follows:
$\rho_{\textrm{f}}=\frac{1}{\dfrac{x_{\textrm{s}}}{\rho_{\textrm{s}}}+\dfrac{1-x_{\textrm{s}}}{\rho_{\textrm{i}}}},$ (A.6)
$C_{p,\textrm{f}}=x_{\textrm{s}}C_{p,\textrm{s}}+(1-x_{\textrm{s}})C_{p,\textrm{i}},$ (A.7)
$k_{\textrm{f}}=x_{\textrm{s}}k_{\textrm{s}}+(1-x_{\textrm{s}})k_{\textrm{i}}.$ (A.8)
Here we assume that the final mass of ice is equal to the initial mass of water (complete freezing), meaning that the amount of water loss due to evaporation during VISF and unfrozen water (bound water) is negligible, which is generally true.
Finally, the height/thickness of the product $H$ can be calculated by
$H=\frac{(m_{\textrm{s}}+m_{\textrm{w,0}})}{\rho_{\textrm{f}}A_{z}}.$ (A.9)
Note that, for some experiments that focus on primary drying, $\rho_{\textrm{f}},C_{p,\textrm{f}},k_{\textrm{f}},$ and $H$ are reported, and so the above calculations may not be needed.
For the secondary drying model (Section 3.4), the volume of the dried region is
$V_{\textrm{e}}=\pi d^{2}H,$ (A.10)
which is constant because there is no moving boundary in this case.
Appendix B Numerical methods
This appendix comprehensively describes numerical techniques required for solving the model equations and summarizes the final equations implemented in MATLAB.
For the freezing model, since a lumped thermal capacity approximation is used, the resulting equations are all ODEs, and so the equations can be solved directly with the selected ODE solver, which is ode15s in this case.
For the primary drying model, the PDE (28) is discretized spatially to produce a system of ODEs, which is then fed to the ODE solver; the technique is referred to as the method of lines [41]. Since the PDE (28) involves a moving interface $S(t)$, define a new variable
$\xi=\frac{z-S}{H-S},$ (B.1)
where $\xi$ represents is dimensionless position with respect to the moving interface (aka sublimation front). This dimensionless position varies from 0 to 1, in which $\xi=0$ at $z=S$ and $\xi=1$ at $z=H$. Consequently, the PDE (28) and boundary conditions (34) and (35) become
$\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{\partial T}{\partial t}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\dfrac{\partial^{2}T}{\partial\xi^{2}}-\frac{\rho_{\textrm{f}}C_{p,\textrm{f}}(\xi-1)}{H-S}\frac{dS}{dt}\frac{\partial T}{\partial\xi}+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\qquad 0<\xi<1,$ (B.2)
$-\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}=h_{\textrm{b}}(T-T_{\textrm{b}}),\qquad\xi=1,$ (B.3)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\dfrac{\partial T}{\partial\xi}+\sigma\varepsilon_{\textrm{gl}}\left(T_{\textrm{c}}^{4}-T^{4}\right),\qquad\xi=0.$ (B.4)
To spatially discretize the PDE, define
$\Delta\xi=\frac{1}{n_{z}-1},$ (B.5)
$\xi=j\Delta\xi,$ (B.6)
where $n_{z}$ is the number of grid points, $\Delta\xi$ is the distance between each grid point, and $j$ is the integer index for discretization, in which $j=1$ at $z=S$ and $j=n_{z}$ at $z=H$. Discretizing (B.2)–(B.4) using the second-order finite difference scheme results in
$\begin{split}\rho_{\textrm{f}}C_{p,\textrm{f}}\dfrac{dT_{j}}{dt}=\frac{k_{\textrm{f}}}{(H-S)^{2}}\!\left(\dfrac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta\xi^{2}}\right)\!-\frac{\rho_{\textrm{f}}C_{p,\textrm{f}}(j\Delta\xi-1)}{H-S}\frac{dS}{dt}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\frac{Q_{\textrm{rad}}}{V_{\textrm{f}}},\\ \qquad j=1,2,\dots,n_{z},\end{split}$ (B.7)
$-\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (B.8)
$N_{\textrm{w}}\Delta H_{\textrm{sub}}=\frac{k_{\textrm{f}}}{H-S}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta\xi}\right)\!+\sigma\varepsilon_{\textrm{gl}}\!\left(T_{\textrm{c}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (B.9)
where $T_{j}$ is the product temperature at position $j$. The ghost point at $j=n_{z}+1$ in (B.8) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (B.7) for $j=n_{z}$. Similarly, $T_{j=0}$ in (B.9) can be eliminated by substituting $T_{j=n_{z}+1}$ into the governing equation (B.7) for $j=1$. The above nondimensionalization and discretization techniques result in a system of nonlinear ODEs defined on a moving-grid system, which is then integrated using ode15s.
Numerical treatment for the secondary drying model are simpler than those described for the primary drying model as there is no moving interface. Define
$\Delta z=\frac{H}{n_{z}-1},$ (B.10)
$z=j\Delta z,$ (B.11)
where $\Delta z$ is the distance between each grid point. Discretizing (38), (39), (41), and (42) using the second-order finite difference scheme results in
$\frac{dT_{j}}{dt}=\frac{k_{\textrm{e}}}{\rho_{\textrm{e}}C_{p,\textrm{e}}}\!\left(\frac{T_{j+1}-2T_{j}+T_{j-1}}{\Delta z^{2}}\right)\!+\frac{\rho_{d}\Delta H_{\textrm{des}}}{\rho_{\textrm{e}}C_{p}}\!\left(\dfrac{dc_{\textrm{w},j}}{dt}\right)\!+\frac{Q_{\textrm{rad}}}{V_{\textrm{e}}},\qquad j=1,2,\dots,n_{z},$ (B.12)
$\dfrac{dc_{\textrm{w},j}}{dt}=f_{\textrm{a}}e^{-E_{\textrm{a}}/RT_{j}}c_{\textrm{w},j},\qquad j=1,2,\dots,n_{z},$ (B.13)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=h_{\textrm{b}}(T_{j}-T_{\textrm{b}}),\qquad j=n_{z},$ (B.14)
$-k_{\textrm{e}}\!\left(\dfrac{T_{j+1}-T_{j-1}}{2\Delta z}\right)\!=\sigma\varepsilon_{\textrm{gl}}\!\left(T_{\textrm{c}}^{4}-T_{j}^{4}\right),\qquad j=1,$ (B.15)
where the ghost points can be treated similarly as done for the primary drying model.
Table 5: *
Nomenclature
Symbols
$A$ area (m2)
$A_{z}$ cross sectional area of the product (m2)
$A_{r}$ side surface area of the product (m2)
$b_{\textrm{n}}$ nucleation kinetics parameter (–)
[…]
$c_{\textrm{w}}$ concentration of bound water (kg water/kg solid)
$c^{*}_{\textrm{w}}$ equilibrium concentration of bound water (kg water/kg solid)
$d$ vial diameter (m)
[…]
$\mathcal{F}$ transfer factor (–)
$F_{1-2}$ view factor from surface 1 to surface 2 (–)
$f_{\textrm{a}}$ frequency factor (1/s)
$H$ height of the product (m)
$\Delta H_{\textrm{des}}$ heat of desorption (J/kg)
$\Delta H_{\textrm{fus}}$ heat of fusion (J/kg)
$\Delta H_{\textrm{sub}}$ heat of sublimation (J/kg)
$\Delta H_{\textrm{vap}}$ heat of vaporization (J/kg)
[…]
$h_{m}$ mass transfer coefficient (kg/m2$\cdots$)
$j$ integer index for discretization (–)
$j_{\textrm{w}}$ mass flow rate of water vapor (kg/s)
$j_{\textrm{w,max}}$ maximum mass flow rate of a condenser (kg/s)
[…]
$k_{\textrm{d}}$ desorption constant (1/s)
[…]
$M$ molar mass (kg/mol)
$m$ mass (kg)
$m_{\textrm{i,n}}$ mass of ice resulting from the first nucleation (kg)
[…]
$n_{\textrm{vial}}$ number of vials (–)
$n_{z}$ number of grid points in the $z$-direction (–)
$P$ probability
$p$ partial pressure (Pa)
$p_{\textrm{t}}$ total pressure (Pa)
$Q$ heat transfer rate (W)
$Q_{\textrm{rad}}$ radiative heat transfer rate (W)
$q$ heat flux (W/m2)
[…]
$R_{\textrm{p}}$ mass transfer resistance of the product during sublimation (m/s)
$R_{\textrm{p0}}$ parameter for the mass transfer resistance (m/s)
$R_{\textrm{p1}}$ parameter for the mass transfer resistance (1/s)
$R_{\textrm{p2}}$ parameter for the mass transfer resistance (1/m)
$r$ radial position (m)
$S$ sublimation front position (m)
$T$ product temperature (K)
$T_{\textrm{b}}$ bottom shelf temperature (K)
$T_{\textrm{c}}$ chamber wall temperature (K)
$T_{\textrm{f}}$ freezing point (K)
$T_{\textrm{g}}$ cryogenic gas temperature (K)
$T_{\textrm{n}}$ nucleation temperature (K)
$\overline{T}$ average chamber temperature (K)
$t$ time (s)
$t_{\textrm{0}}$ initial time, (s)
$t_{\textrm{f1}}$ time when preconditioning completes (s)
$t_{\textrm{f2}}$ time when VISF completes (s)
$t_{\textrm{f3}}$ time when first nucleation completes (s)
$t_{\textrm{f4}}$ time when solidification completes (s)
$t_{\textrm{f5}}$ time when final cooling completes (s)
$t_{\textrm{d1}}$ time when primary drying completes (s)
$t_{\textrm{d2}}$ time when secondary drying completes (s)
$V$ volume (m3)
$x$ mass fraction (–)
$z$ position in vertical direction (m)
$\Delta z$ distance between each grid point in the $z$-direction (m)
$\alpha$ thermal diffusivity, $k/\rho C_{p}$ (m2/s)
$\varepsilon$ emissivity (–)
$\xi$ normalized vertical position (–)
$\Delta\xi$ distance between each grid point in the $\xi$-direction (–)
$\rho$ density (kg/m3)
$\rho_{\textrm{d}}$ density of the dried product (kg/m3)
[…]
Common subscripts
00 initial condition
$1$ surface 1
$2$ surface 2
avg average
c chamber (environment)
b bottom shelf
e effective properties of the dried product (solid + gas)
f frozen material, freezing
gl glass vial
i ice
l liquid solution (water + solute/solid)
n nucleation
s solute/solid
sat saturated condition
s1 top surface of the product
s2 bottom surface of the product
s3 side surface of the product
w water
in inert gas